OpenStudy (anonymous):

Which ordered pair is not a solution of f(x) = 2x + 9? (0, 9) (7, 23) (3, 14) (2, 13)

OpenStudy (leonardo0430):

mother of god 2 rainbow dashes :0 Answr is (2,13)

OpenStudy (anonymous):

haha thnx

geerky42 (geerky42):

You know how ordered pair works? Try input first number in pair, then check if it come out as second number in pair.

OpenStudy (anonymous):

i dont know anything about it, really

geerky42 (geerky42):

None. \((2,13) \) is solution of f(x).

OpenStudy (anonymous):

thnk u both,

geerky42 (geerky42):

\(2(2)+9 = 4+9 = 13\)

OpenStudy (leonardo0430):

yeh no problem :)

OpenStudy (anonymous):

:)

geerky42 (geerky42):

:( lol (2, 13) is not correct answer.

OpenStudy (anonymous):

haha

OpenStudy (leonardo0430):

:0

OpenStudy (anonymous):

lol what is ti?

OpenStudy (anonymous):

it*